# Conghui He4, Xiaoyang Wang6, Fan Wu1

### v2

@@ 2.1 Relational Data Analysis
(2) Agent-Based Methods: Agent-based methods employ multiple LLM agents, each completing tasks under specific prompts, to collaboratively perform analysis. TimeCAP lee2025timecap uses two agents and a multi-modal encoder. The summary agent generates contextual summaries of time series data, while the prediction agent uses these summaries for forecasting. The encoder integrates both data and summaries, reinforced by sampled relevant text, and the final prediction derived from the linear combination of encoder and agent outputs. TimeXL jiang2025explainable extends this with a prototype-based encoder and three agents based on GPT-4o. The encoder produces preliminary predictions and explanations based on time series and textual inputs, while one agent generates predictions from text and explanations, another compares predictions with actual series to provide feedback, and the third refines text quality and triggers encoder updates. The final prediction is a weighted combination of encoder and agent outputs. For context-augmented forecasting wang2024news , a reasoning agent selects relevant news to provide context for predictions, while an evaluation agent critiques both the news selection and prediction against ground truth, updating the reasoning agent’s logic.
@@
(1) Designed Encoder/Decoder: Specialized encoder or decoder modules adapt time series data to LLM input or output formats without modifying the LLM backbone. TIME-LLM DBLP:conf/iclr/0005WMCZSCLLPW24 integrates time series and textual context by encoding time series with a patch reprogram module using text prototypes from pre-trained embeddings, prepending prompts encoded by a frozen LLaMA-7B embedder, and generating forecasts through a trainable projection module. SEED li2025seed bridges structural, numerical, and semantic gaps using a token-aware structural encoder, patch projection and alignment, and semantic reprogramming, concatenating multivariate time series tokens with the task prompt as LLM input. In contrast, TimeCMA liu2025timecma avoids feeding all embeddings into the LLM due to its limitations in learning disentangled representations. It encodes time series through two branches, a time series branch and an LLM-empowered branch, and combines them via cross-modality alignment for multivariate forecasting.
(2) Fine-Tuned LLM Body: Some methods fine-tune LLMs with encoder or decoder modules on time-series tasks, enhancing their understanding and analysis of complex time series. CALF liu2025calf targets embeddings, hidden features, and predictions, applying cross-modal token alignment to derive aligned text tokens queried by time tokens, feature regularization loss between the textual and temporal hidden feature and output consistency loss between textual and temporal prediction to reduce distribution discrepancies during training. S2IP-LLM pan2024sipllm , on the other hand, decomposes time series by seasonal-trend to explicitly encode each component. Tokens of components and relevant prompts are concatenated as input, while the component predictions are combined as the overall prediction. LLM4TS chang2025llm4ts uses a two-stage fine-tuning which aligns LLMs with time series data, then fine-tuning on forecasting tasks. A two-layer aggregation pools the sum of embeddings across time scales to improving temporal understanding ability. Apart from prediction tasks, LLMFew chen2025large addresses multivariate time series classification, using a patch-wise temporal convolution encoder (PTCEnc) to align time series with LLM input, and fine-tunes LLMs via LoRA to enhance feature representation.
@@
TIME-LLM Concatenation Linear Projection
SEED Concatenation Linear Projection
TimeCMA Dual-Modality Multivariate Transformer
CALF Dual-Modality LoRA Dual-Modality
S2IP-LLM Decomposition PEFT Combination
LLM4TS Multi-Scale Two-Stage PEFT RevIN
LLMFew PTCEnc LoRA Classification Head
Relational Data Generation.
On one hand, training or fine-tuning with the training set is essential to adapt LLMs to domain-specific knowledge (e.g., finance, medical, climate) and structured data patterns (e.g., relational, time series, and graph). On the other hand, the test sets are widely used as benchmarks to evaluate the performance of LLMs.
The traditional method of dataset construction is usually the combination of real data collection and manual annotation. For example, the widely used table QA dataset WikiTQ pasupat2015compositional uses tables from Wikipedia and question-label pairs annotated by humans. Similarly, the NL2SQL dataset Spider yu2018spider uses databases collected from existing resources and question-SQL pairs annotated by college students. The correctness, diversity and readability of annotation results are ensured through multiple rounds of manual inspection. In addition, the correctness of SQL is further verified by script. MMTU xing2025mmtu provides a comprehensive collection of 52 such manually curated benchmarks on structured tables, organized across different tasks and benchmark datasets.
However, for some fields, the data may be scattered in small batches, or kept confidential for business or personal privacy needs, making manual collection difficult. To make matters worse, some manual annotation work not only requires a lot of manpower and time costs, but also requires participants to have a high level of professional knowledge. In this situation, semi-automatic or fully automated dataset generation is increasingly needed.
REaLTabFormer solatorio2023realtabformer can generate relational datasets with the transformer. It first generates the parent table with a GPT-2 model and then generates the relational dataset conditioned on that with a Seq2Seq model. The GPT-2 and Seq2Seq model are trained on their corresponding task respectively. Another method hudovernik2024relational utilizes the structural representation capability of GNNs and the generative capability of diffusion models. It represents a relational database by a heterogeneous graph with the method in  xu2022synthetic , where a parent-child table pair connected by a foreign key is modeled as an attributed bipartite graph. Based on the GNN embeddings of the graph, the diffusion-based TabSyn  zhang2023mixed is used for conditional tabular data generation.
@@ 2.2 Graph Data Analysis
Traditional graph analysis (e.g., statistical methods, graph neural network (GNN) based methods) encompasses a spectrum of tasks, including node classification (e.g., categorizing academic papers into research domains), graph classification (e.g., predicting node properties over molecular graphs), link prediction (i.e., inferring latent relationships between graph nodes), community detection (i.e., identifying densely connected subgraphs), anomaly detection (i.e., identifying deviations from expected patterns), graph clustering, and etc. However, these methods have their own limitations. Statistics-based methods fail to handle complex semantic information (e.g., query can be extremely complex and requires human expertise), while graph neural networks (GNNs) exhibit limited generalization capabilities, necessitating task-specific retraining on different tasks.
@@
LLM-Based Semantic Analysis. Furthermore, certain jobs necessitate semantic-aware analysis, such as summarizing textual paragraphs embedded within graph nodes. Based on the adopted LLM strategies, we classify the relevant methods into retrieval-then-reasoning methods, execution-then-reasoning methods, graph-task-based fine-tuning methods, and agent-based methods.
@@
Takeaway: For structured data analysis, LLM-based agents demonstrate significant abilities: Complex Data Understanding (O1): LLMs capture structural and semantic relations across tables, graphs, and time series data, enabling unified understanding. NL-Based Interface (O2): LLMs translate user intents into executable queries (e.g., SQL, graph queries), facilitating intuitive data interaction. Semantic Operators (O3): LLMs integrate symbolic and neural reasoning to execute compositional operations with semantic awareness. Autonomous Evolution (O4): Through continual learning and schema adaptation, LLM agents refine their analytical abilities and evolve with dynamic data and tasks.
@@ 3.1 Markup Language
Figure 4: LLM for Markup Language Overview
@@ 3.2 Semi-Structured Tables
Figure 6: Overview of LLM for Semi-Structured Table.
@@
Semi-Structured Data for LLM To facilitate research on table understanding and reasoning under challenging conditions, a number of datasets have been developed. Below, we introduce several representative datasets that are widely used for semi-structured table understanding.
@@
TEMPTABQA 11–100 1–10 $\times$ $\times$
SPREADSHEETBENCH 101–1K 11–100
MiMoTable 11–100 1–10 $\times$ $\times$
INFOTABS 11–100 1–10 $\times$ $\times$
WikiTableQuestions 101–1K 11–100
FeTaQA 101–1K 1–10 $\times$ $\times$
Takeaway: Semi-structured data remains challenging due to its irregular and nested nature. LLMs show strong potential to address these difficulties through three main directions. Complex Data Understanding (O1): Enable reasoning over irregular and nested structures. NL-based Interfaces (O2): Facilitate intuitive interaction and query generation through natural language. Semantic Operators (O3): Empower models to perform semantic-level operations such as structural interpretation, summarization, and context-aware reasoning.
@@ 4 LLM for Unstructured Data Analysis
Figure 7: Overview of Chart Data Analysis.
@@ 4.1 Chart
Besides, some researches noticed the few-shot ability of LLM and designed relevant end-to-end solutions. ChartGemma masry2024chartgemmavisualinstructiontuningchart , for example, is trained on instruction-tuning data generated directly from chart images, thus exploiting both high-level and low-level information of charts. mPLUG-Owlye2024mplugowlmodularizationempowerslarge is a general-purposed model that also fits the problem. Researchers design a two-stage method to align visual information with text: In the first stage, visual knowledge and abstractor modules are trained; In the second stage, LLM and abstractor modules are fine-tuned using LoRA while freezing visual knowlegde model.
Chart-to-Code. Chart-to-Code refers to generating executable code (e.g., in matplotlib) that can be rendered into chart images, thereby transforming charts into editable forms. Current solutions are primarily end-to-end and rely on VLMs. ChartMimic yang2025chartmimicevaluatinglmmscrossmodal and ChartMoE xu2025chartmoemixturediverselyaligned provide large-scale benchmarks with extensive chart–code pairs to support this task.
@@
Takeaway: Charts present structured data in a visually interpretable form. Future work should focus on mechanisms that emulate human attention for chart understanding, whether through end-to-end or pipeline-based reasoning. Complex Data Understanding (O1): Process charts with different patterns and organize data results mittal-etal-1998-describing ; reiter-2007-architecture ; liu2024chartthinker ; masry2024chartgemmavisualinstructiontuningchart . NL-Based Interface (O2): Pay attention to details that users care about while allowing NL interactions das2025chartsofthoughtenhancingllmvisualization ; ye2024mplugowlmodularizationempowerslarge . Semantic Operators (O3): Capture semantic information hidden in positional relationships among the elements masry2023unichart ; wu2024chartinsights ; xu2025chartmoemixturediverselyaligned . Autonomous Evolution (O4): Notice the distribution of information in charts, and extend its ability to solve a lot of downstream taskshan2023chartllama ; huang2025evochart .
@@ 4.2 Video
Figure 8: Overview of Video Data Analysis.
@@
Takeaway: Through reformulating videos as structured token sequences containing temporal and visual evidence, LLMs can track and understand specific events in object level. Complex Data Understanding (O1): Processing spatial semantics and temporal dynamics, multi-modal cues for emotional analysis, and 3D human behavior wang2024grounded ; weng2025artificial . NL-Based Interface (O2): Driven through natural language using methods including adapters for instruction-conditioned reasoning chen2024timemarker ; li2024llms . Semantic Operators (O3): Perform instruction-aware relative temporal localization to recognize conceptual relationships and contextual details huang2024image ; de2024video . Autonomous Evolution (O4): Advance through adaptive methods like dynamic frame sampling and token merging chen2024timemarker ; yuan2025videorefer .
@@ 4.3 Documents
This section focuses on the challenges and techniques related to visually-rich document understanding, encompassing formats like PDFs, web pages, and scanned reports xu2020layoutlm ; appalaraju2021docformer . Unlike plain text, these documents present information with complex, non-sequential layouts, incorporating visual elements such as tables, lists, figures, and footnotes alongside standard text. The primary challenge lies in parsing this intricate combination of structural, spatial, and semantic information into a coherent, machine-readable representation (e.g., vectors) xu2021layoutlmv2 ; huang2022layoutlmv3 .
@@
Core Challenges in Document Understanding. Document understanding is complicated. For short-form documents (e.g., invoices, web pages), a key bottleneck is form understanding jaume2019funsd , which involves tasks like accurate factual question answering by locating specific information amidst layout noise, generating concise summaries, and performing key-value extraction from structured content. For long-form documents (e.g., scientific papers, technical reports), the challenges escalate to integrative comprehension and cross-document synthesis, along with a robust capability for handling distractors. Finally, handling multilingual corpora requires not only high-fidelity translation but also complex downstream tasks like cross-lingual QA or summarization.
@@
To tackle these challenges, the research community has developed several distinct technical paradigms. We categorize them into three primary approaches: $(i)$ architecture designs ranging from unimodal to multimodal designs, with particular emphasis on fusion strategies and cross-modal interactions, $(ii)$ retrieval-augmented methods for enhancing factual accuracy and context, and $(iii)$ synthetic data generation techniques to address data scarcity and improve model training.
TABLE III: Taxonomy of Multimodal Document Understanding Architectures.
Fusion Stage Architecture Example Characteristics
Early Fusion Single-stream Transformer LayoutLM Concatenated input, unified processing
CNN + Transformer Pipelines Visual CNN, text Transformer, early concat
Intermediate Fusion Dual-stream Transformer LayoutLMv2 DocFormer Parallel text/image streams, cross-attention
Cross-Attention Transformer VLCDoC Inter-modal attention mechanisms
Encoder-Decoder OmniParser CREPE Shared encoder, task-specific decoders
Late Fusion Multi-head Networks DLAFormer Shared backbone, separate output heads
Pipeline Integration OCR+Layout Independent modules, result combination
Unified Processing Vision Transformer ViLT Patch-based, modality-agnostic processing
@@
Architecture Design for Document Understanding. As shown in Table III, the evolution of document understanding architectures reflects a fundamental shift from processing individual modalities in isolation to sophisticated multimodal integration. Early approaches treated text, layout, and visual elements as separate channels, which often led to suboptimal performance due to the loss of critical cross-modal relationships inherent in document structure. Unlike plain text processing, document understanding inherently requires the simultaneous consideration of textual content, spatial layout, and visual appearance. This realization has driven the field toward multimodal architectures that can capture these complex interdependencies.
$\bullet$ Transformer-Based Multimodal Architectures. The transformer architecture has become the dominant paradigm for multimodal document understanding. Foundational models like LayoutLM xu2020layoutlm pioneer the integration of text, layout, and visual information within a unified transformer framework. Building upon this, LayoutLMv2 xu2021layoutlmv2 introduces visual-textual attention mechanisms, while LayoutLMv3 huang2022layoutlmv3 advances unified text–image masking for more effective pre-training. DocFormer appalaraju2021docformer contributes multi-modal self-attention with separate attention heads for different modality pairs. Recent paradigm shifts emphasize efficiency and generalization. For example, DocLLM wang2024docllm represents layout-aware generative models that eliminate the need for raw visual encoders, significantly improving efficiency, whereas ViLT kim2021vilt takes this further by processing visual inputs in a convolution-free manner to concentrate computation on cross-modality interactions.
@@
RAG-Based Document Understanding. To address challenges of factual accuracy and external knowledge, RAG has emerged as a cornerstone architecture (lewis2020retrieval, ). In document understanding, these systems can leverage different evidence modalities.
$\bullet$ Text-Based RAG. The most direct approach involves retrieving from the textual content of documents. More sophisticated methods like RAFT (zhang2024raft, ) focus on optimizing the retrieval and post-processing of text chunks.
$\bullet$ Image-Based RAG. Other systems leverage VLMs to perform retrieval directly on the visual content. M3DOCRAG (M3DOCRAG, ) uses vision-language models to generate embeddings for each document page and computes cross-modal similarity with the query directly on these visual embeddings, completely bypassing OCR. SV-RAG (SV-RAG, ) fine-tunes a VLM to first search for relevant page regions and then generate an answer from the retrieved visual evidence.
$\bullet$ Dual-Modality RAG. Hybrid approaches combine the strengths of both modalities. VisDoM (lei2024visdommultimodal, ), for example, implements parallel visual and textual retrieval branches and synthesizes evidence from both to formulate the final answer, improving reasoning by training models to focus on relevant retrieved context while ignoring distractors (zhang2024raft, ; yan2024corrective, ).

### v3

@@ 2.1 Relational Data Analysis
(2) Step-wise Agent Methods: Step-wise agent methods employ multiple LLM agents, each completing tasks under specific prompts, to collaboratively perform analysis. TimeCAP lee2025timecap employs two agents—a summary agent that generates contextual summaries of time series data and a prediction agent that uses these summaries for forecasting———together with a multi-modal encoder integrating data, summaries, and sampled text.And the final prediction is derived from the linear combination of encoder and agent outputs. TimeXL jiang2025explainable extends this with a prototype-based encoder and three agents. The encoder produces preliminary predictions and explanations based on time series and textual inputs. One agent generates predictions based on textual explanations, another compares them with actual series to provide feedback, and the third refines textual quality and triggers encoder updates. The final prediction is obtained by weighting the outputs of the encoder and agents. For context-augmented forecasting, CABINET wang2024news , a reasoning agent selects relevant news to provide context for predictions, while an evaluation agent critiques both the news selection and prediction against ground truth, updating the reasoning agent’s logic.
@@
(1) Designed Encoder/Decoder: Specialized encoder or decoder modules adapt time series data to LLM input or output formats without modifying the LLM backbone. TIME-LLM DBLP:conf/iclr/0005WMCZSCLLPW24 integrates time series and textual context by encoding time series with a patch reprogram module using text prototypes from pre-trained embeddings, prepending prompts encoded by a frozen LLaMA-7B embedder, and generating forecasts through a trainable projection module. SEED li2025seed bridges structural, numerical and semantic gaps using a token-aware structural encoder, patch projection and alignment, and semantic reprogramming, concatenating multivariate time series tokens with the task prompt as LLM input. In contrast, TimeCMA liu2025timecma avoids feeding all embeddings into the LLM due to its limitations in learning disentangled representations. It encodes time series through two branches, a time series branch and an LLM-empowered branch, and combines them via cross-modality alignment for multivariate forecasting.
(2) LLM Fine-Tuning: Some methods fine-tune LLMs with encoder or decoder modules on time-series tasks, enhancing their understanding and analysis of complex time series. CALF liu2025calf targets embeddings, hidden features, and predictions, applying cross-modal token alignment to derive aligned text tokens queried by time tokens, feature regularization loss between the textual and temporal hidden feature and output consistency loss between textual and temporal prediction to reduce distribution discrepancies during training. S2IP-LLM pan2024sipllm , on the other hand, decomposes time series by seasonal-trend to explicitly encode each component. Tokens of components and relevant prompts are concatenated as input, while the component predictions are combined as the overall prediction. LLM4TS chang2025llm4ts uses a two-stage fine-tuning which aligns LLMs with time series data, then fine-tuning on forecasting tasks. A two-layer aggregation pools the sum of embeddings across time scales to improving temporal understanding ability. Apart from prediction tasks, LLMFew chen2025large addresses multivariate time series classification, using a patch-wise temporal convolution encoder (PTCEnc) to align time series with LLM input, and fine-tunes LLMs via LoRA to enhance feature representation.
@@
TIME-LLM DBLP:conf/iclr/0005WMCZSCLLPW24 Concatenation Linear Projection
SEED li2025seed Concatenation Linear Projection
TimeCMA liu2025timecma Dual-Modality Multivariate Transformer
CALF liu2025calf Dual-Modality LoRA Dual-Modality
S2IP-LLM pan2024sipllm Decomposition PEFT Combination
LLM4TS chang2025llm4ts Multi-Scale Two-Stage PEFT RevIN
LLMFew chen2025large PTCEnc LoRA Classification Head
Relational Data Curation. Training or fine-tuning LLMs on domain-specific datasets is essential for adapting them to specialized knowledge (e.g., finance, medicine, climate) and structured data patterns (e.g., relational, time series, graph). Meanwhile, test datasets serve as benchmarks for evaluating t performance.
The traditional method of dataset construction is usually the combination of real data collection and manual annotation. For example, the widely used table QA dataset WikiTQ pasupat2015compositional uses tables from Wikipedia and question-label pairs annotated by humans. Similarly, the NL2SQL dataset Spider yu2018spider uses databases collected from existing resources and question-SQL pairs annotated by college students. Annotation quality, diversity, and readability are ensured through multiple mannuel reviews, with SQL correctness verified by scripts. MMTU xing2025mmtu provides a comprehensive collection of 52 such manually curated benchmarks on structured tables, organized across different tasks and benchmark datasets.
However, in some domains, data are scattered in small batches or kept confidential for business and privacy reasons, making manual collection difficult. Moreover, manual annotation is not only labor- and time-intensive but also demands substantial domain expertise. In such cases, semi-automatic or fully automatic dataset generation becomes increasingly necessary.
REaLTabFormer solatorio2023realtabformer generates relational datasets with the transformer. It first generates the parent table with a GPT-2 model and then generates the relational dataset conditioned on that with a Seq2Seq model. Another method hudovernik2024relational utilizes the structural representation capability of GNNs and the generative capability of diffusion models. It represents a relational database by a heterogeneous graph with the method in xu2022synthetic , where a parent-child table pair connected by a foreign key is modeled as an attributed bipartite graph.
@@ 2.2 Graph Data Analysis
Traditional graph analysis methods, including statistical approaches and graph neural networks (GNNs), cover a wide range of tasks such as node classification (e.g., categorizing papers by research domain), graph classification (e.g., predicting node properties over molecular graphs), link prediction (i.e., inferring latent relationships), community detection (i.e., identifying densely connected subgraphs), anomaly detection (i.e., identifying deviations from expected patterns), graph clustering, and etc. However, these methods have their own limitations. Statistics-based methods fail to handle complex semantic information, while graph neural networks (GNNs) exhibit limited generalization capabilities, necessitating task-specific retraining on different tasks.
@@
LLM-Based Semantic Analysis. Furthermore, some tasks require semantic-aware analysis, such as summarizing textual content embedded within graph nodes. Depending on the adopted LLM strategies, existing approaches can be categorized into four types: $(i)$ retrieval-then-reasoning methods, $(ii)$ execution-then-reasoning methods, $(iii)$ graph-task-based fine-tuning methods, and (iv)(iv)agent-based methods.
@@
Takeaway (Structured Data): O1: LLMs capture structural and semantic relations across tables, graphs, and time series data, enabling unified understanding. O2: LLMs translate user intents into executable queries (e.g., SQL, graph queries) for intuitive data interaction. O3: LLMs integrate semantic understanding to execute compositional operations with semantic awareness. O4: LLM agents refine the analysis pipeline and evolve with dynamic data and tasks to boost analytical abilities.
@@ 3.1 Markup Language
Figure 4: LLM for Markup Data Analysis.
@@ 3.2 Semi-Structured Tables
Figure 6: LLM for Semi-Structured Table Analysis.
@@
Semi-Structured Data for LLM. To facilitate research on table understanding and reasoning under challenging conditions, a number of datasets have been developed. Below, we introduce several representative datasets that are widely used for semi-structured table understanding.
@@
TEMPTABQA gupta2023temptabqa 11–100 1–10 $\times$ $\times$
SPREADSHEETBENCH ma2024spreadsheet 101–1K 11–100
MiMoTable li2024mimotable 11–100 1–10 $\times$ $\times$
INFOTABS gupta2020infotabs 11–100 1–10 $\times$ $\times$
WikiTableQuestions pasupat2015compositional 101–1K 11–100
FeTaQA nan2022fetaqa 101–1K 1–10 $\times$ $\times$
Takeaway (Semi-Structured Data): O1: LLMs enable reasoning over irregular and nested markup and table structures. O2: LLMs facilitate intuitive interaction and query generation through natural language. O3: LLMs empower semantic-level operations such as structural interpretation, summarization, and context-aware reasoning. O4: Autonomous evolution ability of such analysis needs further research.
@@ 4 LLM for Unstructured Data Analysis
Figure 7: LLM for Chart Data Analysis.
@@ 4.1 Chart
Besides, some researches noticed the few-shot ability of LLM and designed relevant end-to-end solutions. ChartGemma masry2024chartgemmavisualinstructiontuningchart , for example, is trained on instruction-tuning data generated directly from chart images, thus exploiting both high-level and low-level information of charts. mPLUG-Owlye2024mplugowlmodularizationempowerslarge is a general-purposed model that also fits the problem. Researchers design a two-stage method to align visual information with text: In the first stage, visual knowledge and abstractor modules are trained; in the second stage, LLM and abstractor modules are fine-tuned using LoRA while freezing visual knowledge model.
Chart-to-Code. Chart-to-Code refers to generating executable code (e.g., matplotlib) that can be rendered into chart images, thereby transforming charts into editable forms. Current solutions are primarily end-to-end and rely on VLMs. ChartMimic yang2025chartmimicevaluatinglmmscrossmodal and ChartMoE xu2025chartmoemixturediverselyaligned provide large-scale benchmarks with extensive chart–code pairs to support this task.
@@
Takeaway (Chart Analysis): (O1): Process charts with different patterns and organize data results mittal-etal-1998-describing ; reiter-2007-architecture ; liu2024chartthinker ; masry2024chartgemmavisualinstructiontuningchart . (O2): Pay attention to details that users care about while allowing NL interactions das2025chartsofthoughtenhancingllmvisualization ; ye2024mplugowlmodularizationempowerslarge . (O3): Capture semantic information hidden in positional relationships among the elements masry2023unichart ; wu2024chartinsights ; xu2025chartmoemixturediverselyaligned . (O4): Notice the distribution of information in charts, and extend its ability to solve a lot of downstream taskshan2023chartllama ; huang2025evochart .
@@ 4.2 Video
Figure 8: LLM for Video Data Analysis.
@@
Takeaway (Video Analysis): (O1): Processing spatial semantics and temporal dynamics, multi-modal cues for emotional analysis, and 3D human behavior wang2024grounded ; weng2025artificial . (O2): Driven through natural language using methods including adapters for instruction-conditioned reasoning chen2024timemarker ; li2024llms . (O3): Perform instruction-aware relative temporal localization to recognize conceptual relationships and contextual details huang2024image ; de2024video . (O4): Advance through adaptive methods like dynamic frame sampling and token merging chen2024timemarker ; yuan2025videorefer .
@@ 4.3 Documents
Documents present information with complex, non-sequential layouts, incorporating visual elements such as tables, lists, figures, and footnotes alongside standard text (e.g., PDFs, web pages). which involves tasks like accurate factual question answering by locating specific information amidst layout noise, generating concise summaries, and performing key-value extraction from structured content. For long-form documents (e.g., scientific papers, technical reports), the challenges escalate to integrative comprehension and cross-document synthesis, along with a robust capability for handling distractors. Finally, handling multilingual corpora requires not only high-fidelity translation but also complex downstream tasks like cross-lingual QA or summarization.
The primary challenge lies in parsing this intricate combination of structural, spatial, and semantic information into a coherent, machine-readable representation (e.g., vectors) xu2021layoutlmv2 ; huang2022layoutlmv3 .
@@
To tackle these challenges, the research community has developed several distinct technical paradigms. We categorize them into two primary approaches: $(i)$ visual-based analysis, ranging from unimodal to multimodal designs, with particular emphasis on fusion strategies and cross-modal interactions, $(ii)$ text-based analysis for enhancing factual accuracy and context. Finally, we introduce the synthetic data generation techniques to address data scarcity and improve model training for document understanding.
@@
Architecture Design for Document Understanding. The evolution of document understanding architectures reflects a fundamental shift from processing individual modality in isolation to multimodal integration. Early approaches treated text, layout, and visual elements as separate channels, which often led to suboptimal performance due to the loss of critical cross-modal relationships inherent in document structure. Unlike plain text processing, document understanding inherently requires the simultaneous consideration of textual content, spatial layout, and visual appearance. This realization has driven the field toward multimodal architectures that can capture these complex interdependencies.
@@
Visual-Based Document Understanding. The diverse types of elements in documents challenge the accurate document understanding. To solve the challenge, an intuitive way is to apply VLMs to enable visual understanding over such documents.
DocOwl1.5 hu2024mplugdocowl15 is a model for visual document understanding trained on document, chart, and table data. It designs a vision-to-text module which maintains the layout information and reduces the length of visual features through merging horizontal adjacent patches through convolution. Based on DocOwl1.5, DocOwl2 hu2024mplugdocowl2 incorporates a layout-aware compress module to compress each high-resolution document image into a fixed length of 324 tokens, which further improves the efficiency. For multi-page documents, M3DOCRAG (M3DOCRAG, ) uses ColPali faysse2025colpali to embed document pages and the query, and compute cross-modal similarity to retrieve top-k relevant pages for multi-page visual inference. SV-RAG (SV-RAG, ) fine-tunes a VLM to first search for relevant page regions and then generate an answer from the retrieved visual evidence. VisDoM (lei2024visdommultimodal, ) implements parallel visual and textual retrieval branches and synthesizes evidence from both to formulate the final answer, improving reasoning by training models to focus on relevant retrieved context while ignoring distractors (zhang2024raft, ; yan2024corrective, ).
Text-Based Document Understanding. Apart from the visual-based methods, another approach is to convert the document into pure text format and perform text-based inference.
ZenDB lin2024zendb assumes that many documents within a collection share similar templates that encode a common semantic structure. Accordingly, it first performs layout analysis to segment document elements, then applies visual clustering and LLM-based semantic analysis to construct a document tree. Visual elements such as charts and tables are converted into textual descriptions, enabling text-based RAG to retrieve relevant information for question answering. EVAPORATE arora2023language transforms text-based documents into relational data by prompting LLMs to generate table schemas relevant to the query, synthesize functions to retrieve the necessary data, and aggregate the results into a unified relational table, which is then used to answer the original question. Similarly, QUEST sun2025quest builds retrieval indices for both documents and their segments by summarizing and embedding their contents, thereby enabling document querying through SQL-like operations. UQE dai2024uqequery accepts queries in a Universal Query Language (UQL), a dialect of SQL that provides full natural language flexibility in specifying conditions and operators. The UQL is applied on the unstructured text data for data retrieval and question answering. AOP wang2025aop defines semantic operators for building execution workflows (e.g., semantic retrieval, filtering, aggregation). Given an online query, AOP extracts relevant operators and uses these operators to automatically and interactively compose optimized pipelines with the assistance of LLMs to ensure high accuracy question answering. DocETL shankar2025docetl is a document agent which provides a declarative YAML-based interface to define pipelines with LLM-specific operators. It first logically rewrites the analysis pipeline and evaluates the plan guided by a LLM agent. An optimization algorithm is applied to find promising plans, considering the latencies of agent-based plan generation and evaluation.